Consider a classroom that is roughly 5 m X 10m x 3m. Initially t 20oC and P i atm. There are50 people in an insulated class loosing energy to the room at the average rate of 150 watt perperson. How long can they remain in class if the body tcmpcraturc is 37 C and person feels uncomfortable above this temperature. Molar heat capacity of air (72) R.

Grade:12th pass

1 Answers

Arun
3 years ago
Volume of room V=5×10×3×10 
3
  lit
 
P=1atm,R=0.082 atmLmol 
−1
 K 
−1
 ,T=30+273=303K
 
no. of moles of air n= 
RT
PV
​ 
 =6.1×10 
3
 
 
Heat produced per second by '50' students 
=50×150=7500Watt/sec
 
ΔH 
air
​ 
 =nC 
P
​ 
 .ΔT
7500×t×60=6.1×10 
3
 ×7/2×8.314×15
 
t≈5.73 minutes